Find the coordinates of point P that is 1/3 of the way from A(4,-3) to B(1,3). Write your answer as an ordered pair (x, y).
Answer by Alan3354(69443)

Find the coordinates of point P that is 1/3 of the way from A(4,-3) to B(1,3). Write your answer as an ordered pair (x, y).
-----------------
Get the x and y distances separately and add 1/3 to point A:
diffx = -3
1/3 = -1
Px = 4 + (-1) = 3
-------------------
diffy = +6
1/3 = +2
Py = -3 + 2 = -1
-----------------
P is (3,-1)